# MVDiff: Scalable and Flexible Multi-View Diffusion for 3D Object Reconstruction from Single-View

### v1

@@ 2.1 Diffusion for 3D Generation
More recently, SyncDreamer [19] suggests a 3D global feature volume, in order to tackle inconsistencies in multi-view generation. 3D volumes are used with depth-wise attention for maintaining multi-view consistency. The heavy 3D global modeling tend to reduce the speed of the generation and quality of the generated meshes. MVDream [35] on the other hand incorporates 3D self-attention with improved generalisability to unseen datasets.
@@ 4 Experiments
Input condition views are chosen in a similar way as Zero123 [18]. An azimuth angle is randomly chosen from one of the eight discrete angles of the output cameras. The elevation angle is randomly selected in the range [$-10^{\circ},45^{\circ}$]. For data quality purposes, we discard empty rendered images. This represents about one per cent of the training data. 3D objects are centered and we apply uniform scaling in the range [-1,1] so that dimensions matches. Input images to our pipeline are RGB images 256x256.
@@ 4.1 Novel View Synthesis
Zero123 800K 1 18.51 0.856 0.127 7s 12.13 0.601 0.421 7s
Zero123-XL 10M 1 18.93 0.856 0.124 8s 12.61 0.620 0.381 8s
MVDiff 800k 1 20.24 0.884 0.095 9s 12.66 0.638 0.342 9s
MVDiff 800k 2 22.91 0.908 0.064 9s 13.42 0.685 0.321 10s
MVDiff 800k 3 24.09 0.918 0.052 10s 13.58 0.741 0.301 11s
MVDiff 800k 5 25.09 0.927 0.043 11s 14.55 0.833 0.288 12s
MVDiff 800k 10 25.90 0.935 0.036 12s 14.51 0.657 0.215 13s
Table 1: Novel view synthesis performance on GSO and NeRF Synthetic datasets. MVDiff outperforms Zero-123XL with significantly less training data. Additionally, MVDiff performance exhibits further improvement with the inclusion of more reference views.
@@
Point-E 1 0.0561 0.2034
Shape-E 1 0.0681 0.2467
One2345 1 0.0759 0.2969
LGM 1 0.0524 0.3851
SyncDreamer 1 0.0493 0.4581
@@
Table 2: 3D reconstruction performance on GSO dataset. MVDiff outperforms other image-to-3D baselines in generating high-quality 3D objects, with improved performance for multiple input views.
Figure 4: 3D reconstruction from single-view on GSO dataset. MVDiff produces consistent novel views and improves the 3D geometry compared to baselines. One-2-3-45 and SyncDreamer tend to generate overly-smoothed and incomplete 3D objects, in particular the sofa.
@@ 4.3 Ablation Study
Effect of Epipolar and Mult-View Attention.
@@ Acknowledgements
E.B is supported by the Centre for Doctoral Training in Sustainable Approaches to Biomedical Science: Responsible and Reproducible Research (SABS: R3), University of Oxford (EP/S024093/1). P.B. is supported by the UKRI CDT in AI for Healthcare http://ai4health.io (Grant No. P/S023283/1).

### v2

@@ 2.1 Diffusion for 3D Generation
More recently, SyncDreamer [21] suggests a 3D global feature volume, in order to tackle inconsistencies in multi-view generation. 3D volumes are used with depth-wise attention for maintaining multi-view consistency. The heavy 3D global modeling tend to reduce the speed of the generation and quality of the generated meshes. MVDream [37] on the other hand incorporates 3D self-attention with improved generalisability to unseen datasets. EscherNet [18] proposed to leverage camera positional encoding (CaPE) in a transformer-based diffusion model to implicitely learn 3D representations with impressive generalisation and consistency.
@@ 4 Experiments
Input condition views are chosen in a similar way as Zero123 [20]. An azimuth angle is randomly chosen from one of the eight discrete angles of the output cameras. The elevation angle is randomly selected in the range [$-10^{\circ},45^{\circ}$]. For data quality purposes, we discard empty rendered images. This represents about one per cent of the training data. The data filtering strategy is similar to  [18]. 3D objects are centered and we apply uniform scaling in the range [-1,1] so that dimensions matches. Input images to our pipeline are RGB images 256x256.
@@ 4.1 Novel View Synthesis
Zero123[20] 800K 1 18.51 0.856 0.127 7s 12.13 0.601 0.421 7s
Zero123-XL[20] 10M 1 18.93 0.856 0.124 8s 12.61 0.620 0.381 8s
EscherNet[18] 800k 1 20.24 0.884 0.095 8s 13.36 0.659 0.291 9s
MVDiff 800k 1 20.24 0.883 0.094 9s 12.66 0.638 0.342 9s
MVDiff 800k 2 22.92 0.91 0.063 9s 13.42 0.685 0.321 10s
MVDiff 800k 3 24.11 0.921 0.049 10s 13.58 0.741 0.301 11s
MVDiff 800k 5 25.24 0.931 0.040 11s 14.55 0.833 0.288 12s
MVDiff 800k 10 25.94 0.937 0.034 12s 14.51 0.657 0.215 13s
Table 1: Novel view synthesis performance on GSO and NeRF Synthetic datasets. MVDiff outperforms Zero-123 as well as Zero-123XL with significantly less training data. MVDiff shows improved performance with the addition of more reference views.
@@
Point-E[26] 1 0.0561 0.2034
Shape-E[16] 1 0.0681 0.2467
One2345[19] 1 0.0759 0.2969
LGM[41] 1 0.0524 0.3851
SyncDreamer[21] 1 0.0493 0.4581
EscherNet[18] 1 0.0314 0.5974
@@
Table 2: 3D reconstruction performance on GSO dataset. MVDiff surpasses most single-view to 3D benchmarks. Note that the performance improves as the number of input views increases.
Figure 4: 3D reconstruction from single-view on GSO dataset. MVDiff produces consistent novel views and improves the 3D geometry compared to baselines. One-2-3-45 and SyncDreamer tend to generate overly-smoothed and incomplete 3D objects, in particular the sofa. EscherNet recovers more of the finer details, as for the hat.
@@ 4.3 Ablation Study
Effect of Epipolar and Multi-View Attention.
@@ Acknowledgements
E.B is supported by the Centre for Doctoral Training in Sustainable Approaches to Biomedical Science: Responsible and Reproducible Research (SABS: R3), University of Oxford (EP/S024093/1). P.B. is supported by the UKRI CDT in AI for Healthcare http://ai4health.io (Grant No. P/S023283/1). We were inspired by the tables design of Eschernet[18] and we thank the authors for their great work.